OpenStudy (anonymous):

What are the minimum, first quartile, median, third quartile, and maximum of the data set? 9, 20, 4, 18, 4, 18, 20, 9 A. minimum 4; first quartile 10; median 16.25; third quartile 19.5; maximum 20 B. minimum 4; first quartile 6.5; median 13.5; third quartile 19.5; maximum 20 C. minimum 4; first quartile 6.5; median 13.5; third quartile 19; maximum 20 D. minimum 4; first quartile 5.25; median 16.25; third quartile 19; maximum 20

OpenStudy (anonymous):

I'm only confused about the 1st and 3rd quartile

OpenStudy (anonymous):

@amistre64 Can you please explain this for me please?

OpenStudy (amistre64):

whats your median?

OpenStudy (amistre64):

also, put your data in order to help see it better

OpenStudy (amistre64):

quartiles suggest quarters, or divide by 4 take the number of data points you have and divide by 4, then multiply that value by 0,1,2,3,4 to get the quartiles

OpenStudy (amistre64):

IF the quartiles have a decimal, round up IF the quartiles are integers, youll have to average the positions

OpenStudy (anonymous):

My median is 13.5 and when I put them in increasing order I got 4, 4, 9, 9, 18, 18, 20, 20 and I don't understand what you're saying on how to solve the quartiles, I'm lost o.o

OpenStudy (anonymous):

lost please explain

OpenStudy (amistre64):

generally, a quartile is a positional value. you have 8 numbers, or positions to choose from a quartile is a division into quarters, or /4 8/4 = 2 is your basic quartile value. Q0 = 0*2 = 0 Q1 = 1*2 = 2 Q2 = 2*2 = 4 Q3 = 3*2 = 6 Q4 = 4*2 = 8 the values here relate to POSITIONS in the ordered set of data

OpenStudy (amistre64):

the trouble here is that all the values are integers, no decimals so we have to average them out with their next highest position

OpenStudy (amistre64):

Q1 for example is the average of the 2nd and 3rd positions 4, 4, 9, 9, 18, 18, 20, 20 2 3 ^^^^ 4+9 = 13/2 = 6.5 Q2 is the average between the 4th and 5th postiions 4, 4, 9, 9, 18, 18, 20, 20 4 5 ^^^^ 9+18 = 27/2 = 13.5 can you find Q3 for me?

OpenStudy (anonymous):

Uhmmm is it 18 + 20 = 38/2 = 19 ?

OpenStudy (amistre64):

yes :)

OpenStudy (anonymous):

Yayyyyy, thank you lol

OpenStudy (amistre64):

youre welcome, and make sure you pick C ;)

OpenStudy (anonymous):

Lol [x thank you, will do.